"If for a particular genome (X), R = 10 and W = 0.5, what is the relative fitness of the reference genotype (Y)?"
Answer #1

to calculate the relative fitness of each genotype (W) , we calculate it by dividing each genotype survival rate / reproductive rate by highest survival or reproductive rate.

here,

R = 10

W1 = 0.5

0.5 = 10/R2

R2 = 10/0.5 = 20 - genotype with highest reproductive rate

this will be considered as reference genotype and its relative fitness will always be 1 i.e 20/20 = 1.0(W2 )

whereas the relative fitness of 3rd genotype will be 10/20 = 0.5(W3)